# Local stability guarantees for data-driven quadratically nonlinear models

### v1

@@ III.2 Proposed extended trapping SINDy algorithm
$\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (35)
@@
where the matrix $\bm{H}$ defined by (21) can be extracted from the model coefficients. If the regularizer, here the $l_{1}$-norm, is convex then the optimization problem is convex. Sawant et al. leverage this advantage and make the matrix $\bm{A}^{S}$ Hurwitz via a post-processing step, making the optimization problem much easier to solve. However, this a-posteriori fixing of the Hurwitz condition for $\bm{A}^{S}$ has the downside that it may take the solution significantly out of the local minimum found during optimization. For this reason, we combine the two loss functions in Eq. (34) and (35) without the energy-preserving constraint to get,
$\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (36)
@@
This loss function promotes a Hurwitz $\bm{A}^{S}$ matrix while minimizing the destabilizing effects of the cubic contributions in $\dot{v}(\bm{y})$. An estimate of the stability radius of this method is given by Kramer [16]. And to apply this estimate, it is required that the last loss term in Eq. (36) is only motivated in the non-energy-preserving nonlinear setting. Moreover, $\bm{A}^{S}$ is defined by Eq. (15b) instead of Eq. (II.3) in this setting since the energy-preserving constraints do not exist here. In principle, one could instead reformulate the last loss term as an inequality constraint, but we refrain from this strategy because it would involve inequality constraints that are not affine in the model coefficients.
We now additionally promote locally stable models without minimizing the nonlinear effect of the systems. Recall from Thm. 2 that under certain conditions, linear-quadratic models can be locally bounded and exhibit a trapping region even if the hard constraint of the energy-preserving structure is relaxed. For ease of notation, we introduce $Q^{\prime}_{ijk}=Q_{ijk}+Q_{jki}+Q_{kij}$ and $\bm{H}_{0}=\mathcal{T}(\bm{Q}^{\prime})$. The objective function now reads,
$\begin{gathered}\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\left[\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}+\frac{\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})}{\eta}\right]\\ \text{s.t.}\quad\|\bm{Q}^{\prime}\|_{max}=\epsilon_{Q},\end{gathered}$ (37)
@@
$\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (38)
$+\eta^{-1}\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})+\beta^{-1}\|\bm{H}_{0}\|^{2}_{F}\biggr{]}.$
One may notice that according to Thm. 2, it might be better if we directly minimize $\|\bm{H}_{0}\|_{\text{max}}$ rather than $\|\bm{H}_{0}\|_{F}$ to utilize the theorem. However from the perspective of optimization, $\|\bm{H}_{0}\|_{\text{max}}$ is not convex so we choose to minimize $\|\bm{H}_{0}\|_{F}$. Matrix norms are all related by some simple factors so this choice is inconsequential.
@@
Eq. (38) is quite similar to the objective function used in Ouala et al[14]. This optimization can be understood as maximizing the stability radius about the origin at $\bm{m}$ by promoting energy-preserving structure in $\bm{Q}$, instead of maximizing the stability radius about the radius by minimizing the norm of $\bm{Q}$ as in Sawant et al[34], which makes the nonlinearities weak. Considering that there is no “true zero” due to the existence of machine precision in computer arithmetic, even the resulting models from Eq. (34) will not be truly globally bounded numerically.
@@
Notice that we still need $\tilde{\bm{A}}^{S}$ to be Hurwitz to get the estimated stability radius, and the largest eigenvalue $\lambda_{1}$ of $\tilde{\bm{A}}^{S}$ decides the stability radius directly. By additionally optimizing for the Lyapunov matrix $\bm{P}$, as in Goyal [15], we may be able to better control the magnitude of $\lambda_{1}$.
III.3 Interpretation of the loss term $\|\bm{H}_{0}\|^{2}_{F}$
@@
Although the optimization settings of Eq. (36) and Eq. (38) seem to be similar and can both promote models that are locally stable by construction, the results are noticeably different. For Eq. (36), this approach produces stable data-driven models with large stable radii by weakening the nonlinearities to get linearized models. However we could lose much information of the nonlinear behaviors of the physical systems using this optimization, and is marginally helpful for ones who are interested in the behavior of the true models at large distances from the origin. In other words, this method of directly punishing the strength of the nonlinearities requires that the nonlinearities of the systems to be weak so that the data-driven models can perform well in identification tasks.
@@
Instead of minimizing the $\|\bm{H}\|_{F}$, Eq. (38) minimizes $\|\bm{H}_{0}\|_{F}$. To explain further, we do not promote the local stability of the models by weakening the nonlinearities of the system, but by enforcing a weakly quadratic energy-preserving structure. Moreover, we here use this optimization to find the $\rho_{+}$ and $\rho_{-}$, which guarantee the stability of the model based on Thm. 2. Specifically, the size of the domain of boundedness $\rho_{+}$ and the size of the trapping region $\rho_{-}$ is derived from the optimized $\|\bm{H}_{0}\|_{max}$. While Ouala et al [14] make a number of important contributions and extensions to neural-network-based system identification with stability, they are missing the $\rho_{+}$ and $\rho_{-}$ estimates derived in the present work that quantify the stability regions.
@@
We can also reproduce the results from the work by Kaptanoglu et al[13] through Eq. 38, which means our method could also promote global stability in quadratically nonlinear models. As what we show in Sec. II.5, global stability of the models can be achieved, theoretically, by reducing the loss of $\|\bm{H}_{0}\|_{F}$ to machine precision.
@@ IV Results
Table 2: Description of the sampling, extended trapping SINDy hyperparameters, and identified local stability estimates for the dynamic systems examined in Sec. IV. $r$ denotes the dimension of the model. $\beta$ and $\eta$ control the minimization of the two terms in Eq. (38). $\epsilon_{Q}$ reports the deviation from $\bm{Q}$ having zero totally symmetric part. $\lambda_{1}$ denotes the largest eigenvalue in each $\tilde{\bm{A}}^{S}$ of the system.
@@
Lorenz Attractor 3 0.01 $1.78\times 10^{-7}$ $10^{-4}$ $1.3\times 10^{-3}$ $200$ $221$ 8.3 -0.95 1.5 $10^{-5}$
10% noisy Lorenz 3 0.01 $2.27\times 10^{-9}$ $10^{-4}$ $1.1\times 10^{-3}$ 220 243 9.3 -0.92 1.1 0.06
50% noisy Lorenz 3 0.01 $10^{-10}$ $10^{-4}$ $10.0$ 00 00 0.05 1.3 0.9
Finance 3 0.1 $10^{-9}$ $10^{-3}$ $1.1\times 10^{-9}$ 32.7 $5.4\times 10^{7}$ 32.7 -0.1 0.18 $10^{-7}$
Hadley 3 0.015 $10^{-9}$ $10^{-3}$ $3.5\times 10^{-10}$ 22 $1.7\times 10^{8}$ 22 -0.1 0.003 $10^{-9}$
LorenzStenflo 4 0.039 $10^{-9}$ $10^{-3}$ $6.1\times 10^{-7}$ 184 $6\times 10^{4}$ 184 -0.1 0.03 $10^{-7}$
VallisElNino 4 0.022 $10^{-9}$ $10^{-3}$ $3.3\times 10^{-7}$ $3.8\times 10^{3}$ $1.7\times 10^{5}$ $3.8\times 10^{3}$0.1 0.7 $10^{-5}$
We also demonstrate the results qualitatively with dynamical trajectories of the model in order to examine, for instance, the noise robustness of the computational model. Another reason to exhibit results qualitatively rather than quantitatively is that quantifying the model performance for chaotic systems can be challenging because of sensitivity to initial conditions and aperiodic long-term behavior. However, despite the different ways we present for each example, we conduct each experiment in a similar approach. For each system, a single trajectory is used to train a SINDy model with the extended trapping optimizer. Keeping the same temporal duration, we then evaluate the model on different trajectories with new random initial conditions. Table 2 summarizes the sampling, hyperparameters, and identified local stability estimates for a number of examples discussed in Sec. IV.1-IV.2. Several chaotic systems with different strange attractors will be presented in this section. Since it is usually difficult to quantify model performance for a chaotic system, we utilize Eq. (IV) to investigate the utility of our extended trapping algorithm on these systems and compute the time-average error $E_{f}$ in $\dot{\bm{X}}$, rather than calculating the relative prediction error $E_{\text{pred}}$.
@@ IV.1 Noisy Lorenz attractor
The results indicate that the algorithm can improve robustness to noise with respect to the identified model stability. Fig. 5 illustrates that with $10\%$ noise added, the attractor can be well identified. With $50\%$ noise added, the identified model is still stable and represents a different strange attractor, although it fails to report a locally stable region. In contrast, using the traditional SINDy algorithm on Lorenz data with $50\%$ added noise also results in unstable models, even close to the origin.
Another observation is that the $\bm{A}^{S}$ matrix of this $50\%$ noise model is not Hurwitz, which has one and only one positive eigenvalue. This implies that even if the $\bm{A}^{S}$ is not negative-definite, the quadratically nonlinear system may still be locally stable. The reason could be that local stability holds and it is only our stability estimate that has broken down. Or it may not be fully locally stable, but simply have only a very small or degenerate set of initial conditions that can in principle escape to infinity.
@@ IV.2 Assorted locally stable models from the dysts database
The dysts dataset [38] is a collection of more than 100 chaotic systems ideal for benchmarking system identification methods [39]. Four systems with the energy-preserving structure in the quadratic terms are picked for benchmark testing as shown in Fig 6. The datasets of all systems are normalized so we can compare the performance of one set of combinations of hyperparameters on different dynamical systems. The algorithm successfully recognizes the attractors and performs well on promoting stability in modeling these systems. Table 2 also gives the normalized trapping region size of each system.
@@
The Hadley system is an atmospheric convective cell which is also known as Hadley cell or Hadley circulation [41]. It was first envisioned in 1735 to explain the trade winds by Hadley. One way to describe the physical process of Hadley circulation is through this dynamical systems,
@@
where $B=102$, $C=3$ and $p=0$ results in the chaotic dynamics shown in Fig. 5(d).
@@ IV.3 Von Karman vortex street using the enstrophy
The fluid wake behind a bluff body is characterized by a periodic vortex shedding phenomenon known as a von Kármán street. The two-dimensional incompressible flow past a cylinder is a typical example of such behavior, and has been a benchmark problem for Galerkin models for decades. Early Galerkin models of vortex shedding, based on a POD expansion about the mean flow, captured the oscillatory behavior but were structurally unstable [44]. This was later resolved by Noack et al[2], in which an 8-mode POD basis was augmented with a ninth “shift mode” parameterizing the mean flow deformation. This approach was later formalized with a perturbation analysis of the flow at the threshold of bifurcation [45]. The 9-mode quadratic Galerkin model from the Noack et al. work accurately reproduces all of the key physical features of the vortex street. Moreover, in Schlegel and Noack [1] stability of the quadratic model was proven with $m_{9}=m_{\text{shift}}=\epsilon$, $\epsilon>1$, and $m_{i}=0$ for $i=\{1,...,8\}$.
Recent work by Loiseau et al[33, 46, 47] bypassed the Galerkin projection step by using the SINDy algorithm to directly identify the reduced-order dynamics. This approach has been shown to yield compact, accurate models for low-dimensional systems ($r=2$ or $3$), but preserving accuracy and stability for higher-dimensional systems remains challenging. Higher-dimensional regression problems often become ill conditioned; for example, in the cylinder wake example, the higher modes 3-8 are essentially harmonics of the driving modes 1-2, and so it is difficult to distinguish between the various polynomials of these modes during regression. Because these higher harmonics are driven by modes 1-2, the 3D constrained quadratic SINDy model with modes 1-2 plus the shift mode from Loiseau et al[33] already performs well enough to capture the energy evolution with minor overshoot and correct long-time behavior.
In the original trapping SINDy work [13], new, provably globally stable models for the cylinder wake were found. In the present work, reusing the identical dataset described in that work, we extend these results by providing new locally-stable models, including one in which the enstrophy is used as the Lyapunov matrix in Eq. (23a). Both trapping SINDy models show improved accuracy with respect to the POD-Galerkin model in recovering the true limit cycle dynamics in Fig. 7.
@@
Figure 7: Five-dimensional models, integrated from five random initial conditions (color lines), for the von Karman street from the traditional trapping SINDy using the energy as a Lyapunov function, the analytic POD-Galerkin method, and the locally stable trapping SINDy using the enstrophy. The POD modes from the direct numerical simulation data were used for training and are shown in black lines.

### v3

@@ III.2 Proposed extended trapping SINDy algorithm
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (38)
@@
where the matrix $\bm{H}$ defined by (22) can be extracted from the model coefficients. If the regularizer, here the $l_{1}$-norm, is convex then the optimization problem is convex. Sawant et al. leverage this advantage and make the matrix $\tilde{\bm{A}}^{S}$ Hurwitz via a post-processing step, making the optimization problem much easier to solve. However, this a-posteriori fixing of the Hurwitz condition for $\tilde{\bm{A}}^{S}$ has the downside that it may take the solution significantly out of the local minimum found during optimization. For this reason, we combine the two loss functions in Eq. (37) and (38) without the energy-preserving constraint to get,
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (39)
@@
This loss function promotes a Hurwitz $\tilde{\bm{A}}^{S}$ matrix while minimizing the destabilizing effects of the cubic contributions in $\dot{v}(\bm{y})$. Moreover, $\tilde{\bm{A}}^{S}$ is defined by Eq. (16b) instead of Eq. (17) in this setting since the energy-preserving constraints are not prescribed. In principle, one could instead reformulate the last loss term as an inequality constraint, but we refrain from this strategy because it would involve inequality constraints that are not affine in the model coefficients.
We now additionally promote locally stable models without minimizing the nonlinear effect of the systems. Recall from Thm. 2 that under certain conditions, linear-quadratic models can be locally bounded and exhibit a trapping region even if the hard constraint of the energy-preserving structure is relaxed. For ease of notation, we introduce $Q^{\prime}_{ijk}=Q_{ijk}+Q_{jki}+Q_{kij}$, $\bm{H}_{0}=\mathcal{T}(\bm{Q}^{\prime})$ and $\tilde{\bm{H}}_{0}=\bm{P}^{-\frac{1}{2}}\mathcal{T}\bm{H}_{0}(\bm{P}^{-\frac{1}{2}}\otimes\bm{P}^{-\frac{1}{2}})$. The objective function now reads,
$\begin{gathered}\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\left[\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}+\frac{\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})}{\eta}\right]\\ \text{s.t.}\quad\|\tilde{\bm{H}}_{0}\|_{F}=\epsilon_{Q},\end{gathered}$ (40)
@@
$\bm{\xi},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (41)
$+\eta^{-1}\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})+\beta^{-1}\|\tilde{\bm{H}}_{0}\|_{F}^{2}\biggr{]}.$
@@
Eq. (41) is quite similar to the objective function used in Ouala et al[17]. This optimization can be understood as maximizing the stability radius about the origin at $\bm{m}$ by promoting energy-preserving structure in $\bm{Q}$, instead of maximizing the stability radius about the radius by minimizing the norm of $\bm{Q}$ as in Sawant et al[42], which makes the nonlinearities weak. Considering that there is no “true zero” due to the existence of machine precision in computer arithmetic; even the resulting models from Eq. (37) will not be truly globally bounded numerically. A description of each of the hyperparameters $\gamma,\eta$ and $\beta$ is provided in Table 1.
@@
Notice that we still need $\tilde{\bm{A}}^{S}$ to be Hurwitz to get the estimated stability radius, and the largest eigenvalue $\lambda_{1}$ of $\tilde{\bm{A}}^{S}$ determines the stability radius directly. By additionally optimizing for the Lyapunov matrix $\bm{P}$, as in Goyal et al[18], we may be able to better control the magnitude of $\lambda_{1}$.
Table 1: Description of extended trapping SINDy hyperparameters.
$\gamma$ Specifies the strength of sparsity-promotion through the $l_{1}$ regularizer. For simplicity, all results in this paper use $\gamma=0$.
$\eta$ Specifies how strongly to push the algorithm towards models with Hurwitz $\tilde{\bm{A}}^{S}$.
$\beta$ Determines how strongly to penalize nonzero $\|\tilde{\bm{H}}_{0}\|_{F}$.
III.3 Interpretation of the loss term
@@
Although the optimization settings of Eq. (39) and Eq. (41) seem to be similar and can both promote models that are locally stable by construction, the results are noticeably different. For Eq. (39), this approach produces stable data-driven models with large stable radii by weakening the nonlinearities to get linearized models. However, it necessarily throws out information regarding the nonlinear behavior of the physical systems, especially at large distances from the origin. In other words, this method of directly punishing the strength of the nonlinearities is primarily suitable for systems with weak nonlinearities.
@@
Instead of minimizing the $\|\bm{H}\|_{F}$, Eq. (41) minimizes $\|\tilde{\bm{H}}_{0}\|_{F}$. To explain further, we do not promote the local stability of the models by weakening the nonlinearities of the system, but by enforcing a weakly quadratic energy-preserving structure. Moreover, we here use this optimization to find the $\rho_{+}$ and $\rho_{-}$, which guarantee the stability of the model based on Thm. 2. Specifically, the size of the domain of boundedness $\rho_{+}$ and the size of the trapping region $\rho_{-}$ is derived from the optimized $\|\tilde{\bm{H}}_{0}\|_{max}$. While Ouala et al[17] make a number of important contributions and extensions to neural-network-based system identification with stability, they are missing the $\rho_{+}$ and $\rho_{-}$ estimates derived in the present work that quantify the stability regions.
@@
We can also reproduce the results from the work by Kaptanoglu et al[2] through Eq. 41, which means our method could also promote global stability in quadratically nonlinear models. As what we show in Sec. II.5, global stability of the models can be achieved, theoretically, by reducing the loss of $\|\tilde{\bm{H}}_{0}\|_{F}$ to machine precision.
@@ IV Results
We also demonstrate the results qualitatively with dynamical trajectories of the model in order to examine, for instance, the noise robustness of the computational model. Another reason to exhibit results qualitatively rather than quantitatively is that quantifying the model performance for chaotic systems can be challenging because of sensitivity to initial conditions and aperiodic long-term behavior. However, despite the different ways we present for each example, we conduct each experiment in a similar approach. For each system, a single trajectory is used to train a SINDy model with the extended trapping optimizer. Keeping the same temporal duration, we then evaluate the model on different trajectories with new random initial conditions. Table 2 summarizes the sampling, hyperparameters, and identified local stability estimates for a number of examples discussed in Sec. IV.1-IV.2. Several chaotic systems with different strange attractors will be presented in this section. Since it is usually difficult to quantify model performance for a chaotic system, we utilize Eq. (42) to investigate the utility of our extended trapping algorithm on these systems and compute the time-average error $E_{f}$ in $\dot{\bm{X}}$, rather than calculating the relative prediction error $E_{\text{pred}}$.
Table 2: Description of the sampling, extended trapping SINDy hyperparameters, and identified local stability estimates for the dynamic systems examined in Sec. IV. $r$ denotes the dimension of the model. $\beta$ and $\eta$ control the minimization of the two terms in Eq. (41). $\epsilon_{Q}$ reports the deviation from $\tilde{\bm{H}}_{0}$ having zero totally symmetric part. $\lambda_{1}$ denotes the largest eigenvalue in each $\tilde{\bm{A}}^{S}$ of the system.
@@
Lorenz Attractor 3 0.01 $10^{-10}$ $10^{5}$ $8.2\times 10^{-7}$ $103$ $1.8\times 10^{6}$ 4.2 -0.98 0.013 $10^{-5}$
10% noisy Lorenz 3 0.01 $10^{-14}$ $10^{2}$ $6.6\times 10^{-9}$ $132$ $2.2\times 10^{8}$ 5.27 -0.94 0.89 0.18
50% noisy Lorenz 3 0.01 $10^{-10}$ $10^{4}$ $10.0$ 00 00 0.05 1.3 0.9
Finance 3 0.1 $10^{-9}$ $10^{2}$ $8.2\times 10^{-10}$ 32.6 $1.8\times 10^{8}$ 32.6 -0.1 0.18 $10^{-7}$
Hadley 3 0.015 $10^{-9}$ $10^{2}$ $2.8\times 10^{-10}$ 21.3 $5.3\times 10^{8}$ 21.3 -0.1 0.003 $10^{-9}$
LorenzStenflo 4 0.039 $10^{-9}$ $10^{2}$ $5.2\times 10^{-7}$ 181 $2.9\times 10^{5}$ 181 -0.1 0.03 $10^{-7}$
VallisElNino 4 0.022 $10^{-9}$ $10^{2}$ $6.5\times 10^{-9}$ $2.4\times 10^{3}$ $2.3\times 10^{7}$ $2.4\times 10^{3}$0.1 0.7 $10^{-5}$
Von Karman (enstrophy) $5$ $0.1$ $10^{-10}$ $10^{4}$ $5.9\times 10^{-10}$ $4\times 10^{3}$ $2.9\times 10^{8}$ $2.4\times 10^{3}$ $-0.12$
@@ IV.1 Noisy Lorenz attractor
The results indicate that the algorithm can improve robustness to noise with respect to the identified model stability. To control for differential realizations of noise, we made twenty SINDy models from the same Lorenz training trajectory that receives added noise using realizations drawn from the noise distribution. We found that $17/20$ models were identified successfully with local stability guarantees by the extended trapping algorithm in this work. At $50\%$ noise, this number drops to $10/20$ models. For comparison, the traditional SINDy algorithm produces $0/20$ provably stable models for both $25\%$ and $50\%$ noise added. Using SINDy with fairly tight energy-preserving constraints, but no enforcement of the Hurwitz property of $\tilde{\bm{A}}^{S}$, produces $10/20$ provably stable models at $25\%$ noise and $0/20$ at $50\%$ noise.
Fig. 5 illustrates the results with the extended trapping algorithm. With $10\%$ noise added, the attractor can be well identified. With $50\%$ noise added, in this case the identified model is still stable and represents a different strange attractor, but it fails to report a locally stable region. Another observation is that the $\tilde{\bm{A}}^{S}$ matrix of this $50\%$ noise model is not Hurwitz, which has one and only one positive eigenvalue. This implies that even if the $\tilde{\bm{A}}^{S}$ is not negative-definite, the quadratically nonlinear system may still be locally stable. The reason could be that local stability holds and it is only our stability estimate that has broken down. Or it may not be fully locally stable, but simply have only a very small or degenerate set of initial conditions that can in principle escape to infinity. We offer some potential insight into this scenario in Sec. V.
@@ IV.2 Assorted locally stable models from the dysts database
The dysts dataset [45] is a collection of more than 100 chaotic systems ideal for benchmarking system identification methods [46]. Four systems with the energy-preserving structure in the quadratic terms (and effective nonlinearity) are picked for benchmark testing as shown in Fig 6. The datasets of all systems are normalized so we can compare the performance of one set of combinations of hyperparameters on different dynamical systems. The algorithm successfully recognizes the attractors and performs well on promoting stability in modeling these systems. Table 2 also gives the normalized trapping region size of each system.
@@
The Hadley system is an atmospheric convective cell which is also known as Hadley cell or Hadley circulation [48]. It was first envisioned in 1735 to explain the trade winds by Hadley. One way to describe the physical process of Hadley circulation is through this dynamical system,
@@
where $B=102$, $C=3$ and $p=0$ results in the chaotic dynamics shown in Fig. 6(d). For the four systems, we straightforwardly identify models with large locally stable regions with the final $\rho_{-}$ and $\rho_{+}$ listed in Table 2.
@@ IV.3 Von Karman vortex street using the enstrophy
The fluid wake behind a bluff body is characterized by a periodic vortex shedding phenomenon known as a von Kármán street. The two-dimensional incompressible flow past a cylinder is a typical example of such behavior, and has been a benchmark problem for Galerkin models for decades. Early Galerkin models of vortex shedding, based on a POD expansion about the mean flow, captured the oscillatory behavior but were structurally unstable [51]. This was later resolved by Noack et al[3], in which an 8-mode POD basis was augmented with a ninth “shift mode” parameterizing the mean flow deformation. This approach was later formalized with a perturbation analysis of the flow at the threshold of bifurcation [52].
@@
Figure 7: Five-dimensional models, integrated from three random initial conditions (color lines), for the von Karman street from the traditional trapping SINDy using the energy as a Lyapunov function, the analytic POD-Galerkin method, and the locally stable trapping SINDy using the enstrophy. The POD modes from the direct numerical simulation data were used for training and are shown in black lines.
The 9-mode quadratic Galerkin model from the Noack et al. work accurately reproduces all of the key physical features of the vortex street. Moreover, in Schlegel and Noack [1] stability of the quadratic model was proven with $m_{9}=m_{\text{shift}}=\epsilon$, $\epsilon>1$, and $m_{i}=0$ for $i=\{1,...,8\}$.
Recent work by Loiseau et al[41, 53, 54] bypassed the Galerkin projection step by using the SINDy algorithm to directly identify the reduced-order dynamics. This approach has been shown to yield compact, accurate models for low-dimensional systems ($r=2$ or $3$), but preserving accuracy and stability for higher-dimensional systems remains challenging. Higher-dimensional regression problems often become ill-conditioned; for example, in the cylinder wake example, the higher modes 3-8 are essentially harmonics of the driving modes 1-2, and so it is difficult to distinguish between the various polynomials of these modes during regression. Because these higher harmonics are driven by modes 1-2, the 3D constrained quadratic SINDy model with modes 1-2 plus the shift mode from Loiseau et al[41] already performs well enough to capture the energy evolution with minor overshoot and correct long-time behavior.
In the original trapping SINDy work [2], new, provably globally stable models for the cylinder wake were found. In the present work, reusing the identical dataset described in that work, we extend these results by providing new locally-stable models, including one in which the enstrophy
$\bm{P}_{\text{ens}}=\begin{pmatrix}2.69&0.03&0.01&0&-3.17\\ 0.03&2.76&0&0&2.38\\ 0.01&0&10.64&0.08&-0.06\\ 0&0&0.08&10.65&-0.06\\ -3.17&2.38&-0.06&-0.06&3\end{pmatrix}$ (49)
computed using the DNS, is used as the Lyapunov matrix in Eq. (24). Both trapping SINDy models show improved accuracy with respect to the POD-Galerkin model in recovering the true limit cycle dynamics in Fig. 7.
IV.4 Lid-cavity flow
In fluid mechanics, the lid-driven cavity is commonly used as a benchmark for testing numerical methods [55]. A viscous fluid is in a confined geometry of a 2D rectangular container, and its motion is driven by the tangential moving of a bounding wall. The fluid motion is governed by incompressible Navier-Stokes equations. We then apply the model reduction method previously mentioned, resulting in a POD-Galerkin model. In this work, we select the first 6 POD modes to build a data-driven model.
The flow considered in the present work is the incompressible shear-driven cavity flow at $Re=7500$. The computational domain and boundary conditions are the same as in Callaham et al[56]. We also use the same data from direct numerical simulation (DNS) as in the work. Then the extended trapping SINDy algorithm is applied to compare with direct numerical simulation, 6-modes POD-Galerkin model, 16-modes POD-Galerkin model and a naive SINDy model. As shown in Fig. 8, the extended trapping SINDy model outperforms all other presented data-driven models and indicates strong stability compared to all others, even though the $\tilde{\bm{A}}^{S}$ from the extended trapping SINDy is not fully promoted to negative-definite. The $\tilde{\bm{A}}^{S}$ has one positive eigenvalue which is similar as the one in the $50\%$ noisy Lorenz example in Sec. IV.1. Further hyperparameter tuning might facilitate a fully provably stable model but we found empirically this model already works well for the lid cavity flow.